OpenStudy (anonymous):

In right triangle ABC, ∠C is a right angle and sin A = sin B. What is m∠A? A) 30° B) 45° C) 60° D) 75°

OpenStudy (raden):

it means <A = <B A+B+C=180 A+A+90=180 solve for A

OpenStudy (anonymous):

Soo.. Whats the answer?

OpenStudy (raden):

have you tried first ?

OpenStudy (anonymous):

Yesssss

OpenStudy (anonymous):

it means <A = <B A+B+C=180 A+A+90=180 abc will equal 180 because its a triangle A+B+C=180 we already know there is a 90 degree angle so we will put 90 for B A+90+C=180 you need to reach 180 in a right triangle two angles are the same so half of 90 is 45 45+90+45=180 so the answer is 45 degrees